OpenStudy (jb1625):

Evaluate. Let x = 1. Enter the answer in lowest terms.

OpenStudy (alexandervonhumboldt2):

what is the equation?

OpenStudy (jb1625):

That is

OpenStudy (jb1625):

@AlexandervonHumboldt2

OpenStudy (jb1625):

Please help I will give medal to first or best answer

OpenStudy (jb1625):

@pooja195

zepdrix (zepdrix):

\[\large\rm \frac{4\cdot\color{orangered}{x}\div(5-3)}{2\cdot\color{orangered}{x}\cdot5+9\cdot\color{orangered}{x}}\]So replace all of your x's by 1's,\[\large\rm \frac{4\cdot\color{orangered}{1}\div(5-3)}{2\cdot\color{orangered}{1}\cdot5+9\cdot\color{orangered}{1}}\]And simplify. We'll need to be careful with our order of operations I suppose.

OpenStudy (jb1625):

Ok...

zepdrix (zepdrix):

Hmm they made things a little difficult for us since they're using the division symbol AND fraction notation. This is a really strange problem. Let's just deal with the fraction as a very last step. So let's simplify our `denominator`. What do you get? Multiply before adding.

OpenStudy (jb1625):

So you want to do the bottom part first?

zepdrix (zepdrix):

ya :)

OpenStudy (jb1625):

Ok then

OpenStudy (♪chibiterasu):

It doesn't really matter which you do first since you're gonna do the same for both anyway. :p

OpenStudy (jb1625):

19?

zepdrix (zepdrix):

\[\large\rm \frac{4\cdot\color{orangered}{1}\div(5-3)}{19}\]Mmm k great.

OpenStudy (jb1625):

2?

zepdrix (zepdrix):

Good good good :)

OpenStudy (♪chibiterasu):

yeah

zepdrix (zepdrix):

\[\large\rm \frac{2}{19}\]

OpenStudy (jb1625):

:)

zepdrix (zepdrix):

So that would be in lowest terms, good job

OpenStudy (jb1625):

So that is my answer?

zepdrix (zepdrix):

yes

OpenStudy (jb1625):

Thank you!!!